OpenStudy (anonymous):

exponents/ sqaure roots of 4900

OpenStudy (anonymous):

Is this a question?

OpenStudy (anonymous):

Yea \[\sqrt{?}\]4900

OpenStudy (anonymous):

So your question is, "What is the square root of 4900?" sqrt(4900) = sqrt(49*100) = sqrt(49)*sqrt(100) = 7*10 = 70.

OpenStudy (anonymous):

YES thank you